### v1

# Terrier: A Deep Learning Repeat Classifier

Abstract
Repetitive DNA sequences underpin genome architecture and evolutionary processes, yet they remain challenging to classify accurately. Terrier is a deep learning model designed to overcome these challenges by classifying repetitive DNA sequences using a publicly available, curated repeat sequence library trained under the RepeatMasker schema. Existing tools often struggle to classify divergent taxa due to biases in reference libraries, limiting our understanding of repeat evolution and function. Terrier overcomes these challenges by leveraging deep learning for improved accuracy. Trained on RepBase, which includes over 100,000 repeat families—four times more than Dfam—Terrier maps 97.1% of RepBase sequences to RepeatMasker categories, offering the most comprehensive classification system available. When benchmarked against DeepTE, TERL, and TEclass2 in model organisms (rice and fruit flies), Terrier achieved superior accuracy while classifying a broader range of sequences. Further validation in non-model amphibian and flatworm genomes highlights its effectiveness in improving classification in non-model species, facilitating research on repeat-driven evolution, genomic instability, and phenotypic variation.
Keywords Transposable Elements (TEs), Deep Learning, DNA Sequence Classification, Amphibians, Flatworms
1 Introduction
The advent of long-read sequencing has dramatically increased the availability of high-quality reference genomes and thus the ability to detect repeats across a diversity of organisms Osmanski et al. (2023); Rhie et al. (2021). However, repeat classification remains a significant challenge, as most repeat libraries were created from just a few model species such as Drosophila and Arabidopsis Bao et al. (2015). As a result, most repeat classifier tools often fail to classify repeats in divergent taxa Osmanski et al. (2023); Zuo et al. (2023). This limitation hinders the ability to study repeat diversity, evolution, and functional impacts. Given that repeats constitute up to 85% of eukaryotic genomes, they play a crucial role in shaping genome size and structure Wells and Feschotte (2020). While many repeats are neutral Platt et al. (2018), others have been linked to phenotypic traits such as evolutionary rate Platt et al. (2018), coloration Hof et al. (2016); Varga et al. (2020), and fertility Flemr et al. (2013), and some have been implicated in genomic instability and disease Platt et al. (2018); Senft and Macfarlan (2021). Comprehensive classification of repetitive elements is therefore essential for understanding genome evolution and function.
A common software package for identifying repeats is RepeatModeler Flynn et al. (2020), which integrates three de novo repeat discovery tools: RECON Bao and Eddy (2002), RepeatScout Price et al. (2005) and LtrHarvest/Ltr_retriever Ellinghaus et al. (2008). These methods infer repeat boundaries and family relationships using sequence similarity and structural features. However, their dependence on existing reference libraries for classification limits their effectiveness in non-model taxa, where repeat sequences may be highly divergent or underrepresented in curated databases. Machine learning-based classifiers, such as DeepTE and TERL, have improved classification accuracy by learning patterns beyond simple sequence similarity. Yet, their performance remains constrained by the size and diversity of available training datasets, leading to reduced effectiveness when applied to species with distinct repeat landscapes.
Here, we introduce Terrier, a deep learning model designed to improve repeat classification across eukaryotes. Trained using the expansive RepBase library, Terrier enhances classification accuracy by leveraging deep learning for fast and accurate prediction of Transposable Elements (TEs). We validate Terrier’s performance against similar packages in two model organisms: rice (Oryza sativa) and fruit fly (Drosophila melanogaster). We explore its effectiveness in non-model species of amphibians and flatworms. By expanding the scope of repeat classification, Terrier provides a powerful tool for studying repeat diversity, genome evolution, and repeat-driven phenotypic variation.
2 Previous Approaches
2.1 TEclass
TEclass is a software package for classifying TE consensus sequences Abrusán et al. (2009). It represents sequences as frequency vectors of tetramers and pentamers and employs a hierarchical binary classification strategy using support vector machines. First, it determines whether a sequence is a DNA transposon or a retrotransposon; if the latter, it further classifies it as an LTR or non-LTR element, and if non-LTR, it distinguishes between LINEs and SINEs. To account for variability in TE sequence lengths, separate classifiers are trained for different length bins.
TEclass achieved a sensitivity of over 90% for DNA transposons but only 75% when distinguishing between LINEs and SINEs. The authors attribute this decrease to error propagation from earlier steps in the pipeline.
2.2 REPCLASS
REPCLASS is a Perl-based pipeline for classifying TEs Feschotte et al. (2009). It consists of a homology module which uses the input sequence as a query in a TBlastX search using a reference library. A second module looks at the structure of the sequence, while a third module identifies target site duplication (TSD). These various modules allow REPCLASS to predict with more categories than TEclass. It requires the WU-BLAST package, which is no longer maintained.
2.3 PASTEC
PASTEC Hoede et al. (2014) was introduced to classify sequences at order level according to the Wicker hierarchical TE classification system, covering nine autonomous TE orders and three non-autonomous TE types (LARD, TRIM, MITE). It incorporates several methods similar to REPCLASS, including homology-based searches and structural feature detection. Its key innovation is HMM profile detection, which enables classification using conserved functional domains.
PASTEC demonstrated higher classification accuracy compared to existing tools. On RepBase update 15.09 (after removing redundant sequences), it had a misclassification rate of only 15.8%, significantly lower than TEclass (59.7%) and REPCLASS (33.3%). However, it classified only 38.2% of the available sequences, leaving many unclassified due to insufficient evidence.
2.4 DeepTE
DeepTE Yan et al. (2020) is a deep-learning-based tool for transposable element (TE) classification. It represents sequences as k-mer frequency vectors (ranging from trimers to heptamers) and applies a hierarchical classification approach using a convolutional neural network (CNN) with pooling layers. DeepTE trains eight separate models to perform stepwise classifications, distinguishing TEs from non-TEs and further categorizing them into classes, orders, and superfamilies.
Through systematic evaluation of k-mer sizes (k=3 to 7), DeepTE found that heptamers (k=7) provided the best precision overall, though the optimal k-mer size varied across TE groups. Compared to PASTEC, DeepTE achieved higher sensitivity for most TE categories, detecting more true positives, while PASTEC had fewer false positives in some cases. Moreover, DeepTE was over 18 times faster than PASTEC on a GPU-enabled system.
2.5 TERL
TERL da Cruz et al. (2020) is another deep learning model using convolutional neural networks. TERL first represents the DNA sequences using one-hot encoding. These representations are passed through a succession of three convolution and pooling layers and then given to three fully connected layers to make the classification. It is able to classify TEs into nine orders and 29 superfamilies using the Wicker classification system. It achieved a macro mean F1 score of 85.8% for a dataset derived from RepBase 23.10.
TERL was evaluated against PASTEC and TEclass across multiple experiments. It consistently outperformed TEclass in all metrics. Compared to PASTEC, TERL achieved similar performance in several cases, but PASTEC showed higher precision in some categories. However, TERL was orders of magnitude faster than PASTEC, as PASTEC relies on computationally expensive homology-based searches, while TERL uses raw sequence data and CNN-based feature extraction, allowing for efficient GPU computation.
2.6 TEclass2
Bickmann et al. recently developed a deep learning model for classifying repeated sequences called TEclass2 Bickmann et al. (2023). This model uses a sliding window to produce k-mers, then uses a transformer encoder Vaswani et al. (2017) before making the classification. They used the curated and non-curated Dfam 3.7 database with Repbase version 18 as a dataset and used sixteen superfamilies from the Wicker classification system Wicker et al. (2007) as the prediction categories. They chose the uncurated Dfam database because it is vastly larger than the curated version, while acknowledging that it may produce incorrect results due to misclassified repeat families. They achieved an average accuracy of 79% with a macro-averaged accuracy of 72% on the roughly 132,000 sequences in the validation dataset. TEclass2 outperformed TERL and DeepTE in terms of accuracy but typically fewer TE reached the threshold for classification (see Table 2). The software is available but does not include the trained weights. A trained version of the model is available through a web interface.
3 Data
Figure 1: The tree used to classify the repeat families, showing the number of sequences in RepBase mapped to each node.
Figure 2: The relative proportions of each of the repeat categories in the tree.
We used the 29.10 release of the RepBase database for training and cross-validation. This release has just over 100,000 repeat families, approximately four times the number of families as Dfam version 3.8. A challenge with this approach is that RepBase uses a different classification system than RepeatMasker. To ensure compatibility, outputs must be mapped to RepeatMasker categories. Smit and Hubley (2018) produced a version of RepBase aligned with the RepeatMasker classification system, and its metadata is included in RepeatMasker. [^1] https://github.com/rmhubley/RepeatMasker/blob/master/Libraries/RMRBMeta.emblHowever, this edition of RepBase only has 40,000 families, far fewer than than the latest RepBase release. To address this, the Terrier package includes a translation layer that converts RepBase categories to Dfam. This takes 156 RepBase categories and maps them to 9 RepeatMasker Types and 44 SubTypes corresponding to Order and Superfamily classifications to enhance compatibility across annotation tools. This is the largest number of prediction categories among the software packages surveyed above. These categories allow for the translation of 97,529 consensus sequences (97.1%) from RepBase 29.10. The full listing of the number of families in each category is provided in Fig. 1. Fig. 2 shows the relative number of families for each Order and Superfamily.
We divided the families of the dataset into five stratified cross-validation partitions so that each partition would have the same proportion of each Order and Superfamily.
The consensus sequences of the dataset were converted to the SeqBank format [^2] https://github.com/rbturnbull/seqbank for efficient retrieval of sequence data in binary format for training.
The scripts for performing the preprocessing of RepBase are included in the Terrier package with step-by-step instructions included in the documentation.
4 Methods
Figure 3: The Terrier neural network architecture.
To predict the repeat classification from DNA sequences, we employed a neural network architecture based on TorchApp [^3] https://github.com/rbturnbull/torchapp and the DNA sequence classifier Corgi [^4] https://github.com/rbturnbull/corgi. It uses a simple convolutional architecture and was trained to hierarchically predict the repeat Order and Superfamily (Fig. 3). The model takes the DNA sequence and embeds each type of nucleotide base into a vector space of size $e$. These vectors are then processed with a series $n$ layers which perform a convolution with kernel size $k$, followed by a Rectified Linear Unit (ReLU) activation function, a dropout layer with probability $d$ followed by a max pooling operation which reduces the length of the sequence by a factor of 2. The first convolutional layer has $f$ features and this number of features is increased by a growth factor $g$ at each successive layer. The result of these convolutional layers is averaged globally and this is given to a penultimate linear layer of size $p$ with ReLU activation before going to a hierarchical prediction layer provided by the HierarchicalSoftmax package [^5] https://github.com/rbturnbull/hierarchicalsoftmax. This hierarchical prediction layer has outputs for each Order and Superfamily and calculates the loss for the ground truth using cross entropy. The weighting of the Superfamily component of the loss relative to the Order component is controlled by the hyperparameter $\phi$. The hierarchical model outputs Superfamily predictions only if their probability surpasses a user-defined threshold; otherwise, it defaults to the Order level. It can output a CSV with the probabilities of each category as well as the sequences in FASTA format with the classification written after the sequence ID in the header, ready for downstream analysis.
4.1 Training Procedure and Cross Validation
The models were trained for one hundred epochs with a batch size of 32 using the Adam optimization method (Kingma and Ba, 2014). The learning rate was scheduled according to the ‘1cycle’ policy (Smith, 2018) with the peak learning rate set to $10^{-3}$. The hyperparameters were tuned on the first validation partition using the Optuna hyperparameter optimization library (Akiba et al., 2019). The number of features in the first convolutional layer $f$ was scaled to constrain the total number of multiply and accumulate (MACC) operations to approximately $2\times 10^{10}$. The Superfamily accuracy was used as the optimization criterion for twenty runs. The hyperparameters for the optimal training run are displayed in table 1. These were used for training models on the four remaining cross-validation partitions and the Type and Superfamily accuracies are displayed in Fig. 4. This achieved a mean accuracy across the five cross-validation folds at the Order level of 95.2% and at the Superfamily level of 94.0%. The results are comparable across the five cross-validation folds, meaning that the hyperparameter tuning did not overfit to the first validation partition. A confusion matrix for the Order level, found by concatenating the predictions on the five validation sets is displayed in Fig. 5. The final model was trained on the entire training dataset. Steps for reproducing the training of the final model are provided in the documentation.
Hyperparameter Value
Embedding Size $e$ 18
CNN Layer $n$ 4
Layer Growth Factor $g$ 1.96
Dropout $d$ 0.248
Kernel Size $k$ 7
Penultimate Layer Size $p$ 1953
Superfamily loss weighting $\phi$ 1.02
Table 1: Optimized hyperparameters obtained from 20 tuning runs on the first validation partition.
Figure 4: Results for the five cross-validation folds.
Figure 5: The confusion matrix for the concatenated predictions on the five cross-validation sets at the Order level.
5 Test Results
For testing Terrier, we used the rice and fruit fly repeat datasets provided in Bickmann et al. (2023). We compared the performance of Terrier with DeepTE, TERL and TEclass2 using the raw results of these applications on the two datasets, which are provided in the GitHub repository for TEclass2.666https://github.com/IOB-Muenster/TEclass2/blob/main/tests/TEclass2_comparison.ods The tool that we used for this comparison is included in the Terrier application. The TEclass2 results use threshold values of 0.7 and 0.9 for the Superfamily predictions. We present results with the same threshold values for Terrier. The results are presented in Table 2 and plotted in Fig. 6. The confusion matrices for Terrier with a threshold of 0.9 are shown in Fig. 7. The confusion matrices for the other software packages are displayed in the Terrier documentation along with the steps to reproduce these results.
Figure 6: Superfamily classification accuracy versus proportion classified across four software packages. Preferred results are in the top right corner.
Rice Fruit Fly
Order Superfamily Order Superfamily
Software Classified Accuracy Classified Accuracy Classified Accuracy Classified Accuracy
DeepTE 78.7% 52.5% 29.3% 45.5% 87.6% 34.9% 58.2% 31.9%
TERL 97.3% 63.0% 97.3% 46.6% 73.4% 38.8% 66.0% 16.7%
TEclass2 (0.7) 81.3% 86.9% 81.3% 82.0% 64.0% 78.6% 63.9% 67.2%
TEclass2 (0.9) 66.7% 86.0% 66.7% 82.0% 45.3% 86.4% 45.2% 77.7%
Terrier (0.7) 94.7% 94.4% 90.7% 98.5% 80.8% 87.9% 63.6% 91.0%
Terrier (0.9) 90.7% 97.1% 89.3% 98.5% 65.4% 94.5% 53.7% 96.6%
Table 2: Comparison of classification accuracy across test datasets. The highest values per column are bolded.
Figure 7: Confusion matrix for Terrier on the two test datasets using a threshold of 0.9. Predictions at the ‘Superfamily’ level.
5.1 Rice
The Rice dataset contains 75 TE models. Using an NVIDIA H100 GPU, Terrier classified this dataset in a total wall time of 6.3s. With a threshold of 0.7, Terrier classified 71 out of the 75 models to the ‘Order’ level, of which 67 were correctly classified, resulting in an accuracy of 94.37%. This accuracy surpassed the best result from the other packages, with TEclass2 achieving 86.9%. Raising Terrier’s classification threshold to 0.9 improved accuracy to 97.1%, but reduced the number of classified sequences to 68. At the ‘Superfamily’ level, Terrier classified over 90% (68 out of 75) of sequences, with only one incorrect classification, resulting in an accuracy of 98.5%. This was much higher than the best result for the other packages, which was 82% for TEclass2.
5.2 Fruit Fly
The Fruit Fly dataset contains 667 TE models. Terrier classified this dataset in 9.74s, using the same GPU mentioned above. Using a threshold of 0.7, Terrier classified 80.8% of the sequences at the ‘Order’ level, of which 87.9% were correct. This outperformed all other software packages, including TEclass2, which at a threshold of 0.9, achieved an accuracy of 86.4%, but only classified 45.3% of sequences. Using a higher threshold of 0.9 raises the accuracy to 94.5%, but reduces the number of sequences classified to 65.4%. At the ‘Superfamily’ level, using a threshold of 0.7, Terrier predicted 91% of sequences accurately, surpassing TEclass2 at a higher threshold of 0.9, while classifying a far higher proportion of sequences (63.6% compared with 45.2% for TEclass2). When the threshold was raised to 0.9, Terrier’s accuracy increased to 96.6%, but the number of sequences classified decreased to 53.7%.
Figure 8: Computation time for running Terrier, TERL and DeepTE on flatworm and amphibian TE datasets. The filesize refers to the uncompressed FASTA input files in megabytes. A linear trendlines are shown with the equation written on the right.
Figure 9: Comparison between results from RepeatModeler (left) and Terrier (right) experimental data of 8 flatworms and 51 amphibians.
6 Experimental Data
To assess Terrier’s performance, we applied it to experimental data of TE libraries from 8 flatworm species and 51 amphibians.
We present the overall computation time for running Terrier, TERL and DeepTE on these input files. We tested timings on CPUs (eight Intel Xeon Gold 6254) and GPU (NVIDIA A100) (Fig. 8). In all cases, file size was linearly correlated with computational time. Using the CPUs, Terrier scaled at approximately 72s/MB. DeepTE performed faster at around 44s/MB and TERL was faster still at 2.9s/MB. Terrier performed more than 9 times faster with a GPU relative to using the CPUs, scaling at 7.7s/MB. This was faster than DeepTE at 21.3s/MB but still not as fast as TERL which was extremely fast at 0.1s/MB.
In Fig. 9 we show the Order classifications from both RepeatModeler and Terrier at a threshold of 0.7. The RepeatModeler output gives 73.5% as ‘Unknown’ whereas the Terrier classifications are only 34.4% ‘Unknown’.
One of the flatworm species, Schistosoma mansoni, has 21 TE sequences annotated in the NCBI database. Terrier correctly classified 20 of these sequences as LTRs. The remaining non-LTR sequence belongs to the SR2 subfamily and lacks a terminal repeat region. This was classified by Terrier as a DNA transposon with probability 0.72.
7 Conclusion
In this study, we introduce Terrier, a powerful software package designed for the precise classification of repeats from DNA sequences. Terrier distinguishes itself by incorporating a significantly greater number of prediction categories than other comparable methods, providing a finer level of detail for repeat classification. This expanded classification system enables more accurate and comprehensive assessments, especially when dealing with complex, diverse, or poorly characterized genomic regions.
Leveraging deep learning techniques, Terrier not only improves classification accuracy but also delivers exceptionally fast results with minimal computational overhead. Its ability to process large datasets efficiently is particularly evident when running on GPUs, making it an invaluable tool for both large-scale studies and routine use in genomic research.
Compared to other state-of-the-art deep learning classifiers, Terrier demonstrates superior performance in terms of both classification accuracy and the breadth of sequences it can classify. The tool’s high precision, combined with its broad applicability across a range of species and repeat types, positions it as a leading tool for repeat classification. These capabilities make Terrier an essential tool for researchers studying repeat-driven evolution, genomic instability, and other areas where a detailed understanding of repeats is crucial. By advancing the field of repeat classification, Terrier enables more accurate genomic annotations and supports the exploration of complex genomic features in both model and non-model organisms.
8 Data availability
The software is available under an Apache 2.0 Open Source Licence and can be downloaded and installed from GitHub [^7] https://github.com/rbturnbull/terrier. Weights for the model are available with the 0.2.0 release of the software and this is automatically downloaded when first using the model. We also provide a notebook which can be launched in Google Colab to run Terrier on a cloud GPU and to download the results.
The data underlying this article are available in FigShare, at https://doi.org/10.26188/28578992.
9 Competing interests
No competing interest is declared.
10 Author contributions statement
R.T. conceived and wrote the Terrier software, performed the training. R.T., N.Y. and T.K. prepared the training dataset. R.T. and E.T. ran the software packages on the test datasets. N.Y. prepared the flatworm experimental data. T.K. and E.T. prepared the amphibian experimental data. All reviewed the manuscript.
11 Acknowledgments
This research was supported by The University of Melbourne’s Research Computing Services and by Australian Research Council grants FT190100462 and LP200301370 awarded to L.F.S. We acknowledge the help of Priyanka Pillai, Swetha Gopikumar Sreeja and Rafsan Al Mamun.